Tags:
create new tag
view all tags

Shift Log 20120411

AA, LC, LN, PK, SB.

Shift Plan

  • Find OTR.
  • Measure beam size at LWIP.
  • Alter optics to minimise beam size.
  • Minimise background - aim < 0.5V.
  • Alignment.
  • Collisions search.

Shift Log 20120411

17:30

Access to repair polariser cable for OTR.

17:50

Beam on. OTR data taking.

23:20

OTR data taking finishes. Fitted min beam size at LWIP σy = 1.49 +/- 0.06 um.

LW detector on - nearly saturated (1.5V) gives 12000 ADC detector counts with 16dB of attenuation. Background tuning.

ZH3X seems to have a good effect on the background.

Change back the magnets used to reduce the beam size on the OTR screen - (nominal optics gives 100um x 30um) QM14 -43.32A is nominal setting σy 30um QM14 -88.73A gives σy 1.5um

Nominal setting of QM14 gives 100mV background but too large beam size - change QM14 gradually to optimise beam size/background.

QM14 -63.32A σy = 24um

Attenuation reduced to 10dB after detector, 700mV, 9000 counts with this QM14 -73A, QM12 -21A.

QM14 -73.73A, QM12 -21A reduces background to 5000.

Move QM14 to -88.73A, background 7000, σ 1.5um.

Do a final OTR beam size check and detector voltage, then move to timing and collision search.

Detector voltage at this setting 600mV, beam size 1.2um.

02:35

Timing check - laser not locked to e beam. Locked again. Laser/e beam adjusted to be on top of each other for ~ 4.5V.

Wire scanner detector HV turned on.

Searching for e beam chamber vertical position ~ 460 chamber horizontal position ~ -300

CH: -300 CV: 300 - 600 5um steps - nothing
CH: -300 CV: 450 - 500 70 steps - nothing

Steering through the LWIP region with ZH1FF to flatten horizontal orbit in case photons missing detector (new setting 0.231A).

CH: -300 CV: 400 - 600 100 steps - nothing

Set ZH1FF back to -0.269A (original setting) - horizontal angle good but position shifts by ~ 1mm - need to rescan for e beam. ZV1FF also changed from -0.449 to -0.049A.

New position horizontal: 490.

CH: 490 CV: 500 - 700 100 steps - nothing
CH: 490 CV: 700 - 900 100 steps - nothing
CH: 490 CV: 900 - 1100 100 steps - nothing
CH: 490 CV: 300 - 500 100 steps - nothing

Move to new horizontal position: -500.

CH: -500 CV: 100 - 1100 250 steps - nothing

05:30

Try vertical alignment again using sharp edge of screen rather than notch. Turn down wire scanner detector HV to 800V and move screen away from notch to find flat edge, scan screen into beam and look for sharp change in wire scanner detector.

06:30

Checking the timing - laser and e pulses seem overlapped for too large a change of the phase shifter. Try phase scan over region of best spatial alignment.

  • Vertical scan (CH: 490 CV: 450 - 650, 50 steps), starting at phase 4V and repeating in steps of 1V.
  • Vertical scan (CH: 490 CV: 350 - 550, 50 steps), starting at phase 4V and repeating in steps of 1V.

07:25

Collisions! smile

CH: 490, CV: 540, phase 6V. Most sensitive to phase timing.

1st Collisions.jpg

Figure 1 : First collisions 20120412_0714_lws_xy_errbar_plot.pdf.

1stscan.jpg

Figure 2 : First vertical scan showing collisions 20120412_0723_lws_xy_fit_plot.pdf.

Phase scan to optimise timing. New phase: 5.7

phase.jpg

Figure 3 : Phase scan to optimise collisions 20120412_0753_lws_xy_fit_plot.pdf.

  • Vertical scan - signal centred at 535um. Initial fit estimate σy 5.1 +/-0.3um, offset 8848 +/- 18 ADC counts, centre 535.0 +/- 0.3, amplitude 2332 +/- 139.
  • Horizontal scan - signal centred at 305um.

horizontal.jpg

Figure 4 : Horizontal scan 20120412_0733_lws_xy_fit_plot.pdf.

  • Vertical scan - σy 4.1 +/- 0.2, offset 8465 +/- 47, centre 535.9 +/- 0.1, amplitude 3622 +/- 189.

Phase scan again: peak at 5.5V, ~12000 counts.

Vertical steering ZV1FF from -0.049A to -0.069A. Set file set12apr12_0806

Laser telescope actuator 5mm.

08:05

Quad QM14FF scan:

  • -88.73A (CH:305, CV:500 - 580, 80 steps) /dat/raw/20120412_0805_lws.dat σy 4.9 +/- 0.4, offset 7077 +/- 33, centre 534 +/- 0.3, amplitude 2748 +/- 211.

  • Using longer vertical scan (CH: 305, CV: 490 - 580, 90 steps).

  • -90.73A, /dat/raw/20120412_0818_lws.dat
  • -92.73A, /dat/raw/20120412_0834_lws.dat
  • -86.73A, /dat/raw/20120412_0849_lws.dat
  • -84.73A, /dat/raw/20120412_0903_lws.dat
  • -82.73A, /dat/raw/20120412_0916_lws.dat

typicalvertical.jpg

Figure 5 : Typical vertical scan 20120412_0818_lws_xy_fit_plot1.pdf[.

quadscan.jpg

Figure 6 : Quadrupole Scan.

smallestreasonablescan.jpg

Figure 7 : Smallest scan with reasonable data 20120412_0834_lws_xy_fit_plot1.pdf.

  • CV:540, CH:305, 250 pulse log for uncertainties /dat/raw/20120412_0934_lws.dat

10:00

Telescope scan. Lab movers crashed. LN restored control of telescope mover (QM14FF -88.73A).

  • Vertical scan CH:305, CV: 490 - 580, 90 steps, 10 pulses for telescope actuator = 2mm /dat/raw/20120412_1016_lws.dat
  • Vertical scan CH:305, CV: 490 - 580, 90 steps, 5 pulses for telescope actuator = 0mm /dat/raw/20120412_1030_lws.dat
  • Vertical scan CH:305, CV: 490 - 580, 90 steps, 5 pulses for telescope actuator = 7mm /dat/raw/20120412_1040_lws.dat

telescopescan.jpg

Figure 8 : Laser input beam size scan (as function of 2nd lens position.

Topic attachments
I Attachment History Action Size Date Who Comment
PDFpdf 20120412_0714_lws_xy_errbar_plot.pdf r1 manage 111.6 K 16 Apr 2012 - 10:24 LaurieNevay  
PDFpdf 20120412_0723_lws_xy_fit_plot.pdf r1 manage 131.7 K 16 Apr 2012 - 10:24 LaurieNevay  
PDFpdf 20120412_0733_lws_xy_fit_plot.pdf r1 manage 159.7 K 16 Apr 2012 - 10:24 LaurieNevay  
PDFpdf 20120412_0753_lws_xy_fit_plot.pdf r1 manage 104.7 K 16 Apr 2012 - 10:25 LaurieNevay  
PDFpdf 20120412_0818_lws_xy_fit_plot1.pdf r1 manage 136.5 K 16 Apr 2012 - 10:25 LaurieNevay  
PDFpdf 20120412_0818_quad_scan.pdf r1 manage 11.8 K 16 Apr 2012 - 10:25 LaurieNevay  
PDFpdf 20120412_0834_lws_xy_fit_plot1.pdf r1 manage 136.6 K 16 Apr 2012 - 10:25 LaurieNevay  
Edit | Attach | Watch | Print version | History: r8 < r7 < r6 < r5 < r4 | Backlinks | Raw View | Raw edit | More topic actions

Physics WebpagesRHUL WebpagesCampus Connect • Royal Holloway, University of London, Egham, Surrey TW20 0EX; Tel/Fax +44 (0)1784 434455/437520

Topic revision: r8 - 17 Apr 2012 - LaurieNevay

 
This site is powered by the TWiki collaboration platform Powered by PerlCopyright © 2008-2023 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding RHUL Physics Department TWiki? Send feedback